APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday September 22, 2010 (MJD 55462) ---=== OBSERVING TEAM ===--- Steph Snedden Audrey Simmons ---=== OBSERVING PLAN ===--- MARVELS and Flexible Engineering ---=== OBSERVING SUMMARY ===--- We did not open tonight because of lots of rain and humidity. Collected dithered flats on cartridge 13, both with the default exposure time of 2 seconds and then 10 seconds. We learned how to use the wookie with French's help. We replaced a burnt-out FF bulb at the beginning of the night. Steph checked the fiducial tables and created a new version v1_89. We noticed that the LN2 fill lines freeze over and continuously spill nitrogen when they are not locked in tight enough. Tonight we connected the lines after a recent fill when the lines were already frozen and made sure they were tight, but nitrogen still spilled out. We waited a few hours until the lines were back to room temperature and then connected them without any problems. ---=== OBSERVING LOG ===--- Afternoon: ---------- 00:01Z - Tested all actors. All pinged ok. 00:03Z - Spectro biases. 00:32Z - Checked out telescope. Gathered mirror numbers. Mounted cartridge 13. Did not start ventilating due to nearby rain and high humidity. Night: ------ 01:56Z - French taught us how to use the wookie to move cartridges around. 02:05Z - Attempted to started dithered flats on cart 13, but the lower left FF bulb burnt out (when looking at the laptop by the telescope). We replaced it and left the old one on Dan Long's desk. 02:32Z - Dithered flat sequence (2 seconds) 119669-119696. Hartmanns 119697-98. 04:13Z - Dithered flats (10 seconds) 119699-119720. Hartmanns 119721-22. Steph checked in the new version for the fiducial tables. v1_89. Still lots of rain and humidity. 10:27Z - endNight finished. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:34Z 55 55 17 205 (SSW) 38208 - ---=== TELESCOPE STATUS ===--- 09:45Z - Enclosure is on, doors and louvers are closed, and fans are off. Status at: 10:27Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 13 Counterweights at: 275 lbs LN2 autofill systems: Connected and turned on. SP1 LN2 dewar scale: 135.60 lbs, 20.74 psi SP2 LN2 dewar scale: 108.60 lbs, 23.03 psi ---=== SOFTWARE USED ===--- getVersions 2010-09-23 00:00:10 ----STUI software STUI: 1.0b2 2010-07-21 alerts: trunk+svn114742 apo: trunk+svn114701 boss: v2_3_7 boss.daq: 1.2.14 boss.specMech: 3.0.1 gcamera: v1_2_1 guider: v1_2_0_dfc hub: Tag: UNTAGGED_REVISION 502:504M mcp: v6_6_5 mcp.plc: 25 platedb: v1_0_5 sop: v1_1_4 sos: v1_1_4 tcc: 2.13.0 -----Other software sdssProcedure(sdss3): v3_06 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.5100 -5.5100 Axial B 16.0920 16.0940 Axial C 15.9210 15.9220 Trans D 11.0060 11.0100 Lateral E 10.6299 10.6553 Lateral F 12.2428 12.2301 GALILS Commanded: 10600. -250. 5800. 10250. 13050. 12650. Actual: 10600. -250. 5800. 10250. 13053. 12647. SETMIR VALUES PriDesOrient: 0.00 -13.67 10.53 574.80 742.37 PriOrient: 0.00 -13.67 10.53 574.80 742.37 SECONDARY: ---------- Focus: 0.00 Air Temp.: 12.9 Alt.: 30.214384 MIGS TONIGHT NOMINAL Axial A -7.8360 -7.8320 Axial B -0.0410 -0.0410 Axial C 13.8140 13.8180 Trans D 14.1850 14.1840 GALILS Commanded: 1585331. 1566673. 1598874. -1450. -50. Actual: 1585350. 1566695. 1598880. -1462. -69. SETMIR VALUES SecDesOrient: 1257.00 -9.00 -4.00 -58.00 23.67 SecOrient: 1256.99 -9.01 -4.00 -58.08 23.32 ---=== PROBLEMS IN DETAIL ===---